# Rethinking the Sampling Criteria in Reinforcement Learning for LLM Reasoning: A Competence-Difficulty Alignment Perspective

### v2

@@ 4.3 Reinforcement Learning with CDAS
1:Training set $\{(x_{1},y_{1}),(x_{2},y_{2}),\ldots,(x_{n},y_{n})\}$, Model $M_{0}$, total steps $K$, batch size $|B|$
2:Initialize $C_{0}=0$, $t_{j}=0$, $D_{t_{j}}(x_{j})=0\,(j=1,2,\ldots,N)$
3:for $n=1$ to $K$ do
4:     for $j=1$ to $N$ do
5:         $\mathcal{A}(x_{j},M_{n-1})\leftarrow|C_{n-1}-D_{t_{j}}(x)|.$
6:     end for
7:     Sample $B=B^{-}\cup B^{+}$ based on $\mathcal{A}$
8:     for $(x_{j},y_{j})$ in $B$ do  // GRPO Rollout
9:         Compute pass rate $s_{n}(x_{j})$
10:         Compute rewards $r$ for each sampled response
11:         $t_{j}\leftarrow t_{j}+1$
12:         $D_{t_{j}}(x_{j})\leftarrow\frac{t_{j}-1}{t_{j}}\cdot D_{t_{j}-1}(x_{j})+\frac{1}{t_{j}}\cdot(\sigma(C_{n-1}-D_{t_{j}-1}(x_{j}))-s_{n}(x_{j}))$
13:     end for
14:     $C_{n}\leftarrow-\mathbb{E}_{x}[D_{t_{j}}(x)]$
15:     $M_{n+1}\leftarrow$ Update the policy model $M_{i}$ by maximizing the GRPO objective
16:end for
Practically, for each training step, we first measure the alignment between the previous difficulty and model competence, and then obtain the problem set for rollout based on the symmetric sampling above. After one step, the model’s competence and the difficulty of the problems are dynamically updated. The whole process repeats until the training converges, illustrated in algorithm 1. Note that the update of problem difficulty can be viewed as a fixed-point system, formulated as
$\begin{cases}D^{*}(x)=\sigma(C^{*}-D^{*}(x))-S^{*}(x),\ \ x\in\mathbb{X}\\ C^{*}=-\mathbb{E}_{x}[D^{*}(x)]\end{cases}$ (11)
Due to the properties of the sigmoid function which is a contraction mapping, the system will converge to a unique solution. Refer to appendix C for a detailed proof of the convergence.
Note that since $|B|$ is usually much smaller than the size of the training set, performing a full update of problem difficulties at each step will lead to heavy computational overhead. Instead, for problem $x_{j}$, we record the number of times it is sampled as $t_{j}$ and update its difficulty only when it is sampled. Specifically, at the $i$-th step, the problem difficulty can be updated iteratively by
$D_{t_{j}}(x_{j})=\frac{t_{j}-1}{t_{j}}\cdot D_{t_{j}-1}(x_{j})+\frac{1}{t_{j}}\cdot(\sigma(C_{n-1}-D_{t_{j}-1}(x_{j}))-s_{n}(x_{j})).$ (12)
5 Experiments
5.1 Setups
Dataset Following zeng2025simplerl , we utilize the MATH dataset hendrycksmathdataset2021 for RL training, including 7,500 training samples and 4,500 test samples. We reserve its MATH500 subset as the validation set for RL training.
Baselines We compare CDAS against a range of powerful baselines: (1) Random Sampling, which directly samples the problem randomly from the training dataset and can be viewed as plain GRPO training. (2) Curriculum Sampling team2025kimi , which uniformly samples easy problems first, then turns to harder ones. Here, we use the difficulty level tags included in the MATH dataset, ranging from $1$ to $5$, and conduct training from the middle checkpoint in Random Sampling on problems with difficulty level $\geq 4$. (3) Prioritized Sampling, which is adopted in Kimi k1.5team2025kimi that tracks the pass rate $s$ of each problem and then samples problems based on $1-s$.(4) Dynamic Sampling, adopted in DAPOyu2025dapo , which over-samples and filters out problems with the pass rate equal to $1$ or 00.
Training Details We conduct GRPO training experiments on Qwen-2.5-7B yang2024qwen2p5 using the veRL framework sheng2024hybridflowverl . Following the training recipe in zeng2025simplerl , we set the batch size $|B|=1024$, generating $8$ rollouts for each problem with temperature $1.0$ and maximum response tokens $4096$. The training runs for a total $110$ steps, which equals training $10$ epochs on the whole training set. Considering the stability of iterations, we adopt a warm-up strategy in CDAS that ensures problems are sampled randomly within the first $11$ steps to obtain more accurate initial values for problem difficulties. Refer to appendix A for more training details.
Evaluation For a comprehensive evaluation, we select $7$ mathematical reasoning tasks, including AIME 24&25, AMC23, MATH500 hendrycksmathdataset2021 , Minerva Math lewkowycz2022solvingminerva , Olympiad Bench he2024olympiadbench and GSM8K cobbe2021traininggsm8k . Since the number of problems in AIME24&25 and AMC23 is relatively small, we report the Avg@32 metric to reduce randomness. Standard accuracy is reported for the remaining tasks. More details are shown in appendix B.
@@ 5.1 Setups
Table 1: Performance comparison across different sampling methods on various math benchmarks. Metrics are Avg@32 for AIME/AMC and standard accuracy for others. We present the best results in bold and the second with underline.
Steps Method Average Step Time (s) AIME24 (Avg@32) AIME25 (Avg@32) AMC23 (Avg@32) MATH 500 Minerva Math Olympiad Bench GSM8K Avg.
0 Qwen2.5-7B-base - 0.06 0 9.92 47 17.65 14.52 81.5 24.38
55 Random Sampling 1236 10.00 7.29 49.92 74.20 36.76 39.26 91.66 44.73
Curriculum Sampling 1236 10.00 7.29 49.92 74.20 36.76 39.26 91.66 44.73
Prioritized Sampling 1509 11.77 7.08 49.92 74.00 39.34 40.15 92.04 44.90
Dynamic Sampling 2875 12.19 7.92 51.88 75.00 38.97 38.52 91.36 45.12
CDAS (Ours) 1420 11.56 7.71 49.92 76.20 38.97 40.44 92.19 45.28
110 Random Sampling 1245 12.29 6.98 53.98 75.2 37.13 40.00 92.95 45.50
Curriculum Sampling 1375 12.71 7.19 53.98 76.0 38.60 39.56 92.57 45.80
Prioritized Sampling 1670 15.10 9.27 54.61 75.0 37.50 39.56 91.51 46.08
Dynamic Sampling 3745 15.10 9.58 53.44 77.2 39.34 39.56 92.34 46.65
CDAS (Ours) 1608 14.90 11.77 52.03 75.4 40.44 40.89 91.96 46.77
5.2 Main Results
@@
The main results are summarized in table 1, with training curves illustrated in LABEL:fig:train. Besides the final checkpoint, we also report the intermediate checkpoint at the $55$-th step. The accuracy curve on each benchmark can be found in appendix D. From the results, we have the following findings:
@@
CDAS outperforms plain GRPO and other baselines, achieving the highest average accuracy of $45.28\%$ at the 55-th step and $46.77\%$ the 110-th step. We also find that the improvements offered by Curriculum Sampling ($+0.30\%$ at the 110-th step) and Prioritized Sampling ($+0.58\%$ at the 110-th step) over the Random Sampling baseline are limited. It indicates that relying solely on prior difficulty labels or single-step pass rates tends to be suboptimal.
CDAS demonstrates a substantial efficiency advantage over Dynamic Sampling. While Dynamic Sampling achieves performance comparable to CDAS ($46.65\%$ vs. $46.77\%$) by oversampling and excluding problems with pass rates of 1 or 0, it incurs significantly greater computational overhead, requiring $\mathbf{2.33}$X more computation than CDAS. This result clearly highlights the superior efficiency of CDAS to Dynamic Sampling.
CDAS consistently enhances the performance on more challenging benchmarks. On benchmarks like AIME24/25, Minerva Math and Olympiad Bench, CDAS demonstrates remarkable improvements. For instance, CDAS achieves $11.77\%$ accuracy on AIME25 and $40.89\%$ on Olympiad Bench, notably higher than baselines.
The reward curve in CDAS initially grows and then converges to a median value. From LABEL:fig:train, after 30 steps, the amplitude of reward narrows to within the range of $0.4$ to $0.6$ and then converges around $0.5$. This indicates that CDAS effectively selects problems with difficulties appropriate for the model. On the other hand, since Prioritized Sampling tends to select harder problems, its reward demonstrates a continuous downward trend.
An increased response length is not necessary for better performance. Although we observed a rise in response length in CDAS and Curriculum Sampling, the response length in Dynamic Sampling stabilized after about 15 steps, yet the average accuracy on benchmarks continued to grow.
5.3 Ablations
Since CDAS is built on a fixed-point system, the choice of initial values can affect its convergence. We investigate this impact by removing the warmup phase and starting sampling from the first step.
Figure 7: Ablation study of the warm-up phase
As observed in fig. 7, the training curve exhibits significant fluctuations in the early stages of training (the first $20$ steps) without the warmup stage. Specifically, in the first $6$ steps, the removal of the warmup phase leads to a notable decrease in response length along with a rapid increase in reward, suggesting that the model has over-learned from simpler samples. However, as the convergence of the entire fixed-point system is guaranteed, the two training curves gradually overlap, further corroborating the stability of our framework.
Figure 8: Ablation study of symmetric sampling.
As seen in fig. 8, after removing symmetric sampling, CDAS consistently outperforms the Random Sampling baseline. However, there is still an obvious decline in accuracy, especially in the later stages of training (after $100$ steps). We attribute this phenomenon to the accumulated estimation errors in model competence due to the imbalanced sampling. In the early stages of training, imbalanced sampling can cause the model’s competence to be either overestimated or underestimated. This discrepancy from the true competence accumulates as the number of training steps increases and hinders the convergence of CDAS, impairing the performance in the later stages.
6 Analysis and Discussion
@@
In this section, we conduct a comprehensive analysis and discussion of CDAS. We explore the statistical properties of CDAS in terms of sample utility, investigate its advantages over pass rate-based methods, and validate its generalization to code generation tasks along with effectiveness on larger-scale models.
6.1 Utility of the Sampled Problems
@@
From the perspective of the optimization objective of GRPO, the superior performance of Dynamic Sampling can be attributed to its filtering out of samples that do not contribute to model gradientsyu2025dapo (i.e., those with a pass rate of 00 or $1$). Although CDAS does not explicitly constrain the pass rate in problem selection, its alignment-based symmetric sampling inherently mitigates the issue of oversampling the zero-gradient problems to some extent. As illustrated in fig. 9, the proportion of such zero-gradient problems within the batches sampled by CDAS is consistently lower than that of the other baselines, proving that CDAS can effectively improve the utility of sampled problems. We also observe that the proportion of zero-gradient problems in CDAS exhibits a rapid decline during the early stages of training, followed by a slight increase in the later stages. The sharp decrease in the initial phase can be attributed to the swift correction of problem difficulty from its initial values. The modest rise in the later phase is mainly due to the increasing proportion of zero-gradient problems in the whole MATH training set, leading to more problems with a pass rate of 0 sampled in the batch $B^{+}$.
6.2 Problem Difficulty vs. Pass Rate
Since the problem difficulty in CDAS derived from the pass rates, we explore the relationship between them. As illustrated in fig. 10, problem difficulty and pass rate exhibit an overall negative correlation where problems with lower difficulty tend to have higher pass rates. Interestingly, we find that even among problems with the same pass rate, there can still be considerable differences in their estimated difficulties. To further investigate this phenomenon, we randomly selected two problems with a pass rate of 1 at the final sampling step.
@@
As shown in fig. 10, problem A required 25 samplings to reach a pass rate of 1, whereas problem B achieved a pass rate of 1 after only 6 samplings. Despite both having the same final pass rate in rollout, Problem A is noticeably more difficult than Problem B, as indicated by its average accuracy of $32$ inferences is $0.6875$, which is much lower than that of Problem B (Avg@32 = $1.0$). This validates that CDAS, by leveraging historical information, provides a more accurate and robust measure of problem difficulty.
6.3 Generalization
Extension to Code Generations To further investigate the effectiveness of our approach in diverse domains, we apply CDAS to GRPO training on code generation tasks. Specifically, We aggregate open-source data from Appshendrycks2021apps , Tacoli2023taco , and CodeContestsli2022codecontest , and perform GRPO training on Qwen2.5-Coder-7Bhui2024qwen2d5coder for 100 steps. Evaluation results on the LiveCodeBench v5 jain2025livecodebench are shown in LABEL:fig:lcb. Despite the challenges posed by this task for 7B size models, CDAS consistently outperforms the vanilla GRPO baseline in both pass@8 and pass@1 metrics after just 50 steps.
Effectiveness on Larger LLMs In addition to the strong performance observed on Qwen2.5-7B, we further validate the effectiveness of CDAS on larger LLMs. Specifically, we conduct training on Qwen2.5-14Byang2024qwen2p5 with a batch size of 256 for 200 steps, matching the computational budget of our main experiments. The average accuracy of CDAS and Random Sampling is reported in LABEL:fig:14b. We find that CDAS achieves substantial improvements over the Random Sampling baseline by $+1.47\%$, which is even greater than the improvement observed on the 7B model $+1.27\%$, showing effectiveness on larger models.
7 Conclusion

### v3

@@ Introduction
Input: Training set $\{(x_{1},y_{1}),(x_{2},y_{2}),\ldots,(x_{n},y_{n})\}$, Model $M_{0}$
Parameter: Total steps $K$, batch size $|B|$
Output: Updated model $M_{K}$
1: Initialize $C_{0}=0$, $t_{j}=0$, $D_{t_{j}}(x_{j})=0$ for $j=1,2,\ldots,N$
2:for $n=1$ to $K$ do
3:  for $j=1$ to NN do
4:   $\mathcal{A}(x_{j},M_{n-1})\leftarrow|C_{n-1}-D_{t_{j}}(x)|$
5:  end for
6:  Sample $B=B^{-}\cup B^{+}$ based on $\mathcal{A}$       // GRPO Rollout
7:  for $(x_{j},y_{j})$ in $B$ do
8:   Compute pass rate $s_{n}(x_{j})$
9:   Compute rewards $r$ for each sampled response
10:   $t_{j}\leftarrow t_{j}+1$
11:   $D_{t_{j}}(x_{j})\leftarrow\frac{t_{j}-1}{t_{j}}\cdot D_{t_{j}-1}(x_{j})+\frac{1}{t_{j}}\cdot(\sigma(C_{n-1}-D_{t_{j}-1}(x_{j}))-s_{n}(x_{j}))$
12:  end for
13:  $C_{n}\leftarrow-\mathbb{E}_{x}[D_{t_{j}}(x)]$
14:  Update policy model $M_{n+1}$ by maximizing GRPO objective
15:end for
16:return $M_{K}$
Notably, the dynamic and interdependent update process, where competence and difficulty co-evolve, can be rigorously analyzed as a Fixed-Point System. As training progresses and the model’s parameters converge, we expect the system to reach a stable equilibrium state $(C^{*},D^{*},S^{*})$, which satisfies the following conditions:
$\begin{cases}D^{*}(x)=\sigma(C^{*}-D^{*}(x))-S^{*}(x),\ \ \forall x\in\mathcal{D}\\ C^{*}=-\mathbb{E}_{x}[D^{*}(x)]\end{cases}$ (11)
Since the Sigmoid function is a contraction mapping, we can theoretically prove that this system is guaranteed to converge to a unique solution (proof provided in Appendix B). The physical meaning of this equilibrium state provides important insight into the model’s final capabilities:
  • Converged Pass Rate ($S^{*}(x)$): Represents the stable, expected probability of solving problem $x$ once the system stabilizes.
  • Converged Difficulty ($D^{*}(x)$): Measures the gap between problem complexity and model capability. Positive values denote unsolved challenges; negative values denote mastery.
  • Converged Competence ($C^{*}$): Represents the model’s final, overall capability ceiling on the given task, averaged across all problems.
Experiments
Steps Method AIME24 (Avg@32) AIME25 (Avg@32) MATH 500 Minerva Math Olympiad Bench GSM8K Avg.
0 - 0.06 0.00 47.00 17.65 14.52 81.50 24.38
55 Random Sampling 10.00 7.29 74.20 36.76 39.26 91.66 43.20
Curriculum Sampling 10.00 7.29 74.20 36.76 39.26 91.66 43.20
Prioritized Sampling 11.77 7.08 74.00 39.34 40.15 92.04 44.06
Dynamic Sampling 12.19 7.92 75.00 38.97 38.52 91.36 43.99
CDAS (Ours) 11.56 7.71 76.20 38.97 40.44 92.19 44.51
110 Random Sampling 12.29 6.98 75.20 37.13 40.00 92.95 44.09
Curriculum Sampling 12.71 7.19 76.00 38.60 39.56 92.57 44.44
Prioritized Sampling 15.10 9.27 75.00 37.50 39.56 91.51 44.66
Dynamic Sampling 15.10 9.58 77.20 39.34 39.56 92.34 45.52
CDAS (Ours) 14.90 11.77 75.40 40.44 40.89 91.96 45.89
@@
Table 1: Performance comparison across different sampling methods on various math benchmarks. Metrics are Avg@32 for AIME and standard accuracy for others. We present the best results in bold and the second with underline.
Experimental Setup
Dataset Following Zeng et al. (2025), we utilize the MATH dataset (Hendrycks et al. 2021b) for RL training, including 7,500 training samples and 4,500 test samples. We reserve its MATH500 subset as the validation set for RL training.
Baselines We compare CDAS against a range of powerful baselines: (1) Random Sampling, which directly samples the problem randomly from the training dataset and can be viewed as plain GRPO training. (2) Curriculum Sampling (Team et al. 2025), which uniformly samples easy problems first, then turns to harder ones. Here, we use the difficulty level tags included in the MATH dataset, ranging from 1 to 5, and conduct training from the middle checkpoint in Random Sampling on problems with difficulty level $\geq 4$. (3) Prioritized Sampling, which is adopted in Kimi k1.5 (Team et al. 2025) that tracks the pass rate $s$ of each problem and then samples problems based on 1s1-s.(4) Dynamic Sampling, adopted in DAPO (Yu et al. 2025), which over-samples and filters out problems with the pass rate equal to 1 or 0.
Training Details We experiment with GRPO, the most widely-used algorithm for LLM RL training, on Qwen2.5-7B (Yang et al. 2024) using the veRL framework (Sheng et al. 2024). Following the training recipe in SimpleRL-Zoo (Zeng et al. 2025), we set the batch size $|B|=1024$, generating $8$ rollouts for each problem with temperature $1.0$ and maximum response tokens $4096$. Refer to Appendix C for more training details.
Considering the stability of our iterative system, we adopt a warm-up strategy in CDAS. For the duration of the first training epoch, we sample problems uniformly at random. This ensures every problem is sampled at least once, allowing us to establish a robust initial value for each problem’s difficulty. By doing so, we directly mitigate the risk of amplifying dataset biases before the main adaptive sampling phase begins. Further analysis is available in Appendix F.
Evaluation For a comprehensive evaluation, we select 6 mathematical reasoning tasks, including AIME 24&25, MATH500 (Hendrycks et al. 2021b), Minerva Math (Lewkowycz et al. 2022), Olympiad Bench (He et al. 2024) and GSM8K (Cobbe et al. 2021). Since the number of problems in AIME24 and AIME25 is relatively small, we report the Avg@32 metric to reduce randomness. Standard accuracy is reported for the remaining tasks. More details are shown in Appendix D.
Main Results
@@
The main results are summarized in Table 1, with training curves shown in Figure 4 and performance comparison curves shown in Figure 2. Besides the final checkpoint, we also report the intermediate checkpoint at the 55-th step. From the results, we have the following findings:
@@
CDAS outperforms plain GRPO and other baselines. As shown in Table 1, CDAS achieves the best average accuracy at both the 55-th step (44.51%) and the 110-th step (45.89%) checkpoints. Meanwhile, we find that the improvements offered by Curriculum Sampling (+0.35% at the 110-th step) and Prioritized Sampling (+0.57% at the 110-th step) over the Random Sampling baseline are limited. This indicates that relying solely on prior difficulty labels or single-step pass rates tends to be suboptimal.
CDAS demonstrates a substantial efficiency advantage. In terms of sample efficiency, CDAS achieves a performance of 44.51% at just 55 training steps, surpassing the final performance of Random Sampling (44.09% at 110 steps), reaching comparable capability in half the training time. Moreover, as shown in Figure 2(b), the training time of CDAS is comparable to most standard baselines. This stands in stark contrast to Dynamic Sampling, which incurs significantly higher computational costs (2.33× slower).
The reward curve in CDAS demonstrates a dynamic process. As shown in Figure 4, the reward curve for CDAS exhibits four distinct phases: (1) Warm-up Phase: During the first epoch, the curve mirrors that of Random Sampling, as problems are sampled uniformly to establish an initial, unbiased difficulty assessment. (2) Rapid Ascent Phase: Immediately following the warm-up, the reward sharply increases to over 0.8. This is because CDAS identifies and prioritizes the problems that have become easy for the model. (3) Adaptive Descent Phase: As the model’s competence grows, CDAS intentionally transitions to sampling more challenging problems that better match the model’s new capabilities. This strategic shift causes the average reward to decrease. (4) Convergence Phase: The curve finally stabilizes around a median reward of approximately 0.55, which signifies that CDAS continuously presents the model with problems within its competence.
(a) Random Sampling and Curriculum Sampling
(b) Prioritized Sampling
(c) Dynamic Sampling
(d) CDAS
Figure 4: Training curves of different sampling strategies.
Analysis and Discussion
@@
In this section, we conduct a comprehensive analysis and discussion of CDAS. We explore the statistical properties of CDAS in terms of sample utility, investigate its advantages over pass rate-based methods, and validate its generalization to code generation tasks along with different architectures and model sizes.
Utility of the Sampled Problems
@@
From the perspective of the optimization objective of GRPO, the superior performance of Dynamic Sampling can be attributed to its filtering out of samples that do not contribute to model gradients (Yu et al. 2025) (i.e., those with a pass rate of 0 or $1$). Although CDAS does not explicitly constrain the pass rate in problem selection, its alignment-based symmetric sampling inherently mitigates the issue of oversampling the zero-gradient problems to some extent. As illustrated in Figure 5, the proportion of such zero-gradient problems within the batches sampled by CDAS is consistently lower than that of the other baselines, proving that CDAS can effectively improve the utility of sampled problems. We also observe that the proportion of zero-gradient problems in CDAS exhibits a rapid decline during the early stages of training, followed by a slight increase in the later stages. The sharp decrease in the initial phase can be attributed to the swift correction of problem difficulty from its initial values. The modest rise in the later phase is mainly due to the increasing proportion of zero-gradient problems in the whole MATH training set, leading to more problems with a pass rate of 0 sampled in the batch $B^{+}$.
Problem Difficulty vs. Pass Rate
Since the problem difficulty in CDAS derived from the pass rates, we explore the relationship between them. As shown in Figure 6, problem difficulty and pass rate exhibit an overall negative correlation where problems with lower difficulty tend to have higher pass rates. Interestingly, we find that even among problems with the same pass rate, there can still be considerable differences in their estimated difficulties. To further investigate this phenomenon, we randomly selected two problems with a pass rate of 1 at the final sampling step.
@@
As shown in Figure 6, problem A required 25 samplings to reach a pass rate of 1, whereas problem B achieved a pass rate of 1 after only 6 samplings. Despite both having the same final pass rate in rollout, Problem A is noticeably more difficult than Problem B, as indicated by its average accuracy of 32 inferences being 0.6875, which is much lower than that of Problem B (Avg@32 = 1.0). This validates that CDAS, by leveraging historical information, provides a more accurate and robust measure of problem difficulty.
Generalization
We further investigate the generalization capabilities of CDAS across different tasks, model architectures, and model sizes. We compare our method against two representative sampling strategies: Random Sampling (RS), the most widely used baseline, and Dynamic Sampling (DS), a strong but computationally expensive performance baseline.
Generalization to Code Generation Tasks We apply CDAS to the code generation task. Specifically, we aggregate 10k open-source competitive programming problems from Apps (Hendrycks et al. 2021a), Taco (Li et al. 2023), and CodeContests (Li et al. 2022), and perform GRPO training on Qwen2.5-Coder-7B (Hui et al. 2024) for 100 steps. As shown in Table 2, CDAS achieves 18.24% Pass@1 accuracy on LiveCodeBench v5 (Jain et al. 2025), outperforming RS by +3.00% and nearly matching DS. This verifies CDAS’s strong generalization to different tasks. We leave the comparison of other tasks for future work.
Method Pass@1 Acc.
RS 15.24
DS 18.32
CDAS 18.24
Table 2: Accuracy comparison on LiveCodeBench v5.
Generalization to Different Architectures and Model Sizes To further examine the generalization capability of CDAS across architectures and scales, we conduct additional RL experiments on Qwen2.5-14B (Yang et al. 2024) and OctoThinker-8B-Hybrid (Wang et al. 2025). Both models are trained on the MATH dataset used in the main experiments. For Qwen2.5-14B, we train for 200 steps with a batch size of 256, while OctoThinker-8B-Hybrid adopts the same hyperparameter configuration as the main setup. As shown in Table 3, CDAS consistently outperforms the RS baseline and achieves performance that matches or exceeds the strong DS baseline on both models, while maintaining comparable training efficiency to RS. These results demonstrate the robust generalization of CDAS to models of larger scale and different architectural designs.
Model Method Average Acc.
Qwen2.5-14B RS 46.28
DS 46.62
CDAS 47.22
OctoThinker-8B RS 35.30
DS 36.47
CDAS 36.52
Table 3: Generalization Performance Across Different Architectures and Model Sizes.
Conclusion